OpenStudy (anonymous):

I got x = -2.61 and 4.61 for the proporttion, x/x+ 12 = 1/x -1. Am I correct? I have no idea how to check.

OpenStudy (anonymous):

\[\frac{ x }{ x }+12=\frac{ 1 }{ x }-1\] Is that what your problem looked like? and you said that you got \[x=-2.61\] and this for the proportion? \[4.61\]

OpenStudy (anonymous):

To check your answer you will plug in -2.61 or 4.61 for all x's in the equation

OpenStudy (anonymous):

Yes, and also I feel that you might have messed up some where. Can you show what you did to figure out the answer?

OpenStudy (anonymous):

I cross multiplied x/x + 12 and 1/x-1

OpenStudy (anonymous):



OpenStudy (anonymous):

Look at it as \[13=\frac{ 1 }{ x }-1\] then reverse the inequality in order to isolate the x on the left hand side. \[13=\frac{ 1 }{ x }-1\rightarrow \frac{ 1 }{ x }-1=13\] Isolate the terms with x on the left side \[\frac{ 1 }{ x }=14\] and you should get \[x=\frac{ 1 }{ 14 }\]

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

no, i am still confused

OpenStudy (anonymous):

Hmm. @Notamathgenius help this user so they're no longer confused .-.

OpenStudy (notamathgenius):

Plug in the -2.61 for x

OpenStudy (anonymous):

\[\frac{ 1 }{ -2.16 -1 }\frac{ -2.16 }{ -2.16 + 12 } \]

OpenStudy (anonymous):

like so?

OpenStudy (notamathgenius):

Yes